### v3

# Semi-Dirac fermions in hexagonal close-packed cadmium

Abstract
Semi-Dirac fermions are massless in one direction and massive in the perpendicular directions. Such quasiparticles have been proposed in various contexts in condensed matter. Using first principles calculations, we identify a pair of semi-Dirac bands in the electronic structure of hexagonal close-packed (hcp) cadmium, which keep their linear out-of-plane and quadratic inplane dispersions up to the Fermi level. We demonstrate that the dichotomy between the linear and quadratic dispersions is driven by an orientation-sensitive hybridization between the $s$ and $p_{z}$ orbitals. The upper semi-Dirac band produces a lens-shaped nonellipsoidal Fermi sheet whose cross-section area has a $k$-dependence that is in excellent agreement with the experimentally measured period of Sondheimer oscillations.
Linearly dispersive electronic bands can give rise to massless analogues of relativistic particles in condensed matter systems. Such Dirac fermions lead to intriguing experimental consequences like the $E_{N}\propto\sqrt{NB}$ energy spectrum of the Landau levels, half-integer anomalous quantum Hall effect, and Klein tunneling [1]. Observation of these phenomena in monolayer graphene [2], which has Dirac cones at its Brillouin zone corners, has motivated the search for other such systems.
Semi-Dirac fermions exhibiting direction-dependent massless and massive behaviors have attracted much attention during the last fifteen years [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. They were proposed to exist in the $A$ phase of 3He [3], in a graphenelike model subjected to magnetic field [4], VO2 layers confined within TiO2 [5], strained silicene oxide [14], and tetragonal perovskite oxides with the $I4/mcm$ space group [16]. While their observation have been reported in a microwave analogue of a graphenelike lattice [12], honeycomb lattice of micropillars [15], and in cold atoms [10, 11], they have been more elusive in solids. Recently, a semi-Dirac dispersion leading to a $[(N+\frac{1}{2})B]^{2/3}$ scaling of the Landau levels [4, 6] has been reported in ZrSiS [17].
Generally, linear dispersion is expected when two nonbonding bands cross at a Dirac point situated away from the Brillouin zone center. Semi-Dirac bands in (TiO2)5/(VO2)3 arise in this way [5]. Another avenue is the inversion of the relative positions of bands associated with $s$ and $p$ orbitals. In HgTe-CdTe quantum wells, such an inversion occurs at a critical thickness leading to a Dirac point at the Brillouin zone center [18]. Given that $s$-$p$ hybridization is uniform along all three axes in cubic zinc blende HgTe, however, a semi-Dirac analogue of this mechanism has been missing.
A possible avenue to obtain semi-Dirac bands is to explore cases with direction-dependent $s$-$p$ hybridization. Metals in column twelve of the periodic table are promising, as their mobile electrons originate from $s$ and $p$ orbitals rather than $d$ orbitals. Furthermore, unlike column eleven elements (the noble metals) with a nominal electronic configuration $s^{1}$, column twelve metals are nominally $s^{2}$. The additional electron in column 12 metals favors the hcp structure, where the $p_{z}$ orbital hybridizes differently with the $s$ orbital compared to the $p_{x}$ and $p_{y}$ orbitals.
In this paper, we use first principles calculations to show that hcp Cd hosts a pair of semi-Dirac bands. The bands’ dispersion is parabolic in plane and linear out of the plane over an extended distance in the reciprocal space that includes the Brillouin zone center and the Fermi level crossing. The two bands have mixed $s$ and $p_{z}$ character in the out-of-plane direction and unmixed $s$ or $p_{z}$ character in the in-plane directions. Thus, the direction-dependent mixing of the $s$ and $p_{z}$ characters is responsible for the presence of the semi-Dirac bands. Comparison with hypothetical hcp Ag shows that the relative positions of the $s$- and $p_{z}$-like bands control the direction-dependent mixing. We show that in such a hypothetical hcp metal with a single $s$ electron, there is a large gap between $s$-like and $p_{z}$-like states. Consequently, $s$-$p_{z}$ hybridization is negligible and the bands are parabolic in all directions. We find that tensile strain moves the $s$-like states higher relative to the $p_{z}$ states, thereby increasing the $s$-$p_{z}$ hybridization in the out-of-plane direction. Semi-Dirac dispersion occurs near the point of $s$-$p_{z}$ inversion, in a manner analogous to that of HgTe-CdTe quantum wells where the band inversion is controlled by tuning the strength of the effective spin-orbit interaction. The Fermi level crossing of the upper semi-Dirac band generates a non-ellipsoid Fermi pocket, called lens-shaped in the scientific literature. We find that the calculated derivative of the cross-section area of this sheet with respect to the out-of-plane wave vector agrees with the experimental value extracted from the observed Sondheimer oscillations, corroborating the presence of semi-Dirac fermions in this material.
Figure 1: a) Calculated LDA band structure of hcp Cd. The band structure features a pair of bands that disperse linearly along the out-of-plane direction $\Gamma A$ and quadratically along the inplane directions $\Gamma M$ and $\Gamma K$. These bands appear to converge near $-3$ eV at $\Gamma$, but they are in fact separated by 50 meV. The upper and lower semi-Dirac bands are shown in blue and red, respectively. Calculated Fermi surface of hcp Cd consisting of b) hole-type “cap”, c) hole-type “monster”, and d) electron-type “lens” sheets. The lens sheet derives from the upper semi-Dirac band.
The calculated LDA band structure of hcp Cd (see the supplement [19] for details) is shown in Fig. 1. There are ten narrow bands with $4d$ character between $-10$ and $-8$ eV relative to the Fermi level. Dispersive bands with $5s$ and $5p$ characters lie above and cross the Fermi level. The calculated Fermi surface of hcp Cd is shown in Fig. 1b-d. It consists of two hole-type sheets known as “caps” and “monster” centered around the six edges of the Brillouin zone and the electron-type lens sheet centered at $\Gamma$. The volume enclosed by these sheets are $0.01$ and $0.19$ holes and $0.20$ electrons per cell, respectively. The corresponding carrier concentrations are 0.17, 3.31 and $-3.48\times 10^{22}$ cm-3.
[…]
Figure 2: Calculated LDA band structure of a hypothetical hcp Ag using the lattice parameters of hcp Cd. The pair of bands that linearly disperse along $\Gamma A$ in hcp Cd disperse quadratically in hypothetical hcp Ag. The gap at $\Gamma$ between these bands also increases to 1.4 eV.
Figure 3: Band structures of hcp Cd and hypothetical hcp Ag isostructural to hcp Cd as a function of respective $s$ and $p_{z}$ orbital characters along out-of-plane $\Gamma A$ and inplane $\Gamma M$ and $\Gamma K$ directions. The semi-Dirac bands in hcp Cd show a mixed $s$ and $p_{z}$ character in the out-of-plane direction. In hcp Ag, the bands disperse quadratically in all directions, and the upper band is predominantly $s$-like, while the lower band is mainly $p_{z}$-like.
Semi-Dirac points previously discussed in the literature are associated with degeneracies due to band crossings [4, 5, 6, 9]. In the present case, there is a gap of 50 meV at $\Gamma$ near $-3$ eV where two semi-Dirac bands converge even without spin-orbit interaction. To disentangle the role of structure and chemistry behind the origin of these semi-Dirac bands, let us suppose that the next-door element in the periodic table Ag, instead of being face-centered cubic (its stable structure), could be synthesized in an hcp structure. Fig. 2 shows the band structure of such a hypothetical hcp Ag with the same structural parameters as that of hcp Cd. We can see that the band structures of isostructural hcp Ag and Cd differ by more than a rigid shift of the Fermi level even though they lie next to each other in the periodic table and their electron count differs only by one. Hcp Ag no longer exhibits the linearly dispersing bands along the out-of-plane direction $\Gamma A$. The gap between these bands increases to 1.4 eV, and they taper off parabolically as they approach $\Gamma$. Therefore, a crossing of nonbonding bands do not seem to play a role in the appearance of the semi-Dirac bands in hcp Cd. Furthermore, the contrasting electronic structures of isostructural hcp Cd and Ag shows that the semi-Dirac bands do not arise merely due to the layered hexagonal structure and points out the essential role played by the chemistry of divalent $5s^{2}$ electronic configuration in Cd.
We have plotted the band structures of hcp Cd and Ag showing the contribution of $s$ and $p_{z}$ orbital characters along the out-of-plane and two in-plane directions in Fig. 3 to understand the changes in band hybridization due to the presence of an additional $5s$ electron in Cd. Along the out-of-plane $\Gamma A$ direction, we can see that the pair of semi-Dirac bands in hcp Cd show nearly equal amount of $s$ and $p_{z}$ characters, especially near $\Gamma$. In hcp Ag, in contrast, the upper-lying band shows dominant $s$ character, while the $p_{z}$ character is dominant in the lower-lying band. This is particularly the case close to $\Gamma$, where the upper-lying and lower-lying bands are almost solely due to the $s$ and $p_{z}$ states, respectively. Thus, $s$-$p_{z}$ hybridization along one orientation gives rise to a linear dispersion, and the lack of it along the two others leads to a quadratic one. This observation is further reinforced looking at the orbital characters of these bands along $\Gamma M$ and $\Gamma K$. In both materials, along these in-plane directions, each band exhibits one dominant orbital character, and all of them disperse quadratically.
Interestingly, the higher-lying band along the in-plane directions has $p_{z}$ character in hcp Cd but $s$ character in hcp Ag. This band inversion is reminiscent of the one noted in HgTe-CdTe quantum wells [18]. When the well is thin, the $s$-derived band lies higher than the $p$-derived band, the orbital mixing is low, and the bands are quadratic with a finite gap at $\Gamma$. With increasing thickness of HgTe in the quantum well, the orbital characters start to mix, the gap decreases, and the bands start to become linearly dispersive. The gap vanishes and the bands cross each other linearly at a critical thickness, beyond which the higher lying band exhibits more $p$ character. In this analogy, hcp Ag is like thin HgTe below the critical thickness and its $s$-type band lies above its $p_{z}$-type band. Meanwhile, hcp Cd is analogous to thick HgTe above the critical thickness, because its $p_{z}$-type band lies just above its $s$-type band.
Figure 4: a) The electron-like Fermi surface generated by the upper semi-Dirac band projected in the $k_{x},k_{z}$ plane (empty circles). The blue solid line represents a regular ellipsoid whose axes are equally long. Note the difference. b) The cross section area as a function of $k_{z}$. Over a long distance, near the poles (that is $k_{z,min}$ and $k_{z,max}$), the area is linear in $k_{z}$, consistent with the linear dispersion of the semi-Dirac band from which this Fermi sheet derives.
Since the band structures of hcp Cd and hcp Ag have been calculated using the same lattice parameters, the difference in the relative position of their $s$- and $p_{z}$-type bands near the Fermi level is due to the presence of an additional electron in Cd. The extra electron does not just rigidly fill up the unoccupied states in the band structure when going from hcp Ag to hcp Cd, it also pulls down the $s$-type states relative to the position of the $p_{z}$-type states. We discuss the inversion of the relative positions of the $s$- and $p_{z}$-type states in hcp Ag due to tensile strain in the supplement [19].
Direct observation of the linear dispersion of these semi-Dirac bands in hcp Cd emerges as a task for future experiments employing angle-resolved photoemeission spectroscopy. However, it has already a striking experimental signature, detectable by scrutinizing data published decades ago, reporting robust Sondheimer oscillations [22] with a specific periodicity in cadmium [23].
In a metallic crystal hosting ballistic electrons, galvonometric coefficients display oscillations, which are periodic in magnetic field. The period of these oscillations is given by [24, 25]:
$\Delta B=\frac{\hbar}{ed}\left(\frac{\partial A}{\partial k_{z}}\right),$ (1)
Here, $d$ is the thickness in the direction perpendicular to the magnetic field, $e$ is the electron charge, $\hbar$ is the reduced Planck constant, $k_{z}$ is the component of the wave vector along the magnetic field, and $A$ is the cross section area of the Fermi sheet intersecting with a plane perpendicular to $k_{z}$.
In a perfect ellipsoid, the relevant $\frac{\partial A}{\partial k_{z}}$ is the one at the limit point (the apex) of the ellipsoid, since the degeneracy there is slightly larger than everywhere else. With increasing number of states sharing an identical $\frac{\partial A}{\partial k_{z}}$, oscillations sharing the same periodicity become more robust. In our case, the electron-like pocket formed by the upper semi-Dirac band (Fig.1d) is not an ellipsoid. Fig. 4a shows how the projection of this Fermi surface in the ($k_{x}$, $k_{z}$) plane is not a regular ellipsoid. Fig. 4b shows its cross section area as a function of $k_{z}$. Along this orientation, the apices of this lens-like Fermi surface occur at ±$\pm 0.28$-1. The area is proportional to $k_{z}$ along a large fraction of the out-of-plane path in the Brillouin zone near these points, as expected by the linear dispersion in this direction. The figure also shows a linear fit, which demonstrates that the quadratic component is negligible up to a distance of $0.09$-1 towards the Brillouin zone center from the apices. The slope of the linear fit is $\frac{\partial A}{\partial k_{z}}=8.73$-1. Grenier et al. found a period of $\delta B=0.0564\pm 3T$ in a sample with $d=1.02$ mm. Inserting these in Eq. 1, yields $\frac{\partial A}{\partial k_{z}}=8.76$-1. Such an agreement, better than one percent, has been confirmed by a recent experimental quantifying the periodicity across a forty-fold variation of thickness [26].
In summary, revisiting the electronic structure of hcp Cd using first principles calculations demonstrates that it hosts semi-Dirac bands with linear out-of-plane dispersion and quadratic in-plane dispersion. We argue that this is driven by direction-dependent mixing of $s$ and $p_{z}$ states, and that hcp Cd can be likened to HgTe-CdTe near its critical thickness, except with the restriction of band hybridization along a single direction. The nonellipsoid electron pocket produced by the upper semi-Dirac band has a cross-section area proportional to the out-of-plane wave vector along a significant fraction of the Brillouin zone path and is in excellent quantitative agreement with the period of Sondheimer oscillations.
This work was supported by GENCI-TGCC under grant no. A0150913028.